OpenStudy (anonymous):

. Spherical Wedge. A portion of spherical cheese (queso de bola) was sliced in the form of a spherical wedge. The diameter of the cheese (wax included) is 3 inches. Calculate the volume of the wedge if the angle is 45o. Calculate the outer surface area of the wax taken-off.

OpenStudy (anonymous):

Yes. The one i did was the total surface area

OpenStudy (anonymous):

Help please @paki

OpenStudy (anonymous):

@One098 @KamiBug help

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iambatman

OpenStudy (perl):

we can use this http://en.wikipedia.org/wiki/Spherical_wedge

OpenStudy (anonymous):

The volume i got is 101.25 @perl

OpenStudy (perl):

one moment

OpenStudy (perl):

definition spherical wedge (geometry) A volume of a sphere lying between two vertical planes through the centre of the sphere.

OpenStudy (perl):

in radians there is a nice formula for volume V = 2/3 * a * r^2 , where 'a' is the center angle in radians

OpenStudy (perl):

45 degrees = pi/4 radians V = 2/3 * pi/4 * (3/2)^2 ,

OpenStudy (perl):

inches cubed

OpenStudy (anonymous):

Its (3/2)^3

OpenStudy (perl):

the diameter was given as 3 inches, so r = 3/2

OpenStudy (perl):

oh sorry

OpenStudy (perl):

V = 2/3 * pi/4 * (3/2)^3 *

OpenStudy (perl):

V = 9/16 * Pi

ganeshie8 (ganeshie8):

you may simply scale the volume for the given angle : \[V = \dfrac{\pi/4}{2\pi}\left( \dfrac{4}{3}\pi r^3 \right) = \frac{1}{8}\left( \dfrac{4}{3}\pi \left(\frac{3}{2}\right)^3 \right) = \frac{9\pi}{16}\]

OpenStudy (perl):

right thats basically what you are doing volume of wedge = theta/ 2pi * ( volume of sphere)

ganeshie8 (ganeshie8):

you can do the same thing for surface area also just keep in mind you got two more faces when you chop off a slice

OpenStudy (perl):

and similarly for surface area Surface area of wedge = theta/ 2pi * (surface area of sphere)

OpenStudy (anonymous):

How about the area ?

OpenStudy (anonymous):

3pi/4 b

OpenStudy (perl):

similarly for outer surface area of wedge Outer Surface area of wedge = theta/ 2pi * (surface area of sphere) area = (pi/4) / (2pi ) * 4 * pi * r^2

OpenStudy (anonymous):

?*

OpenStudy (perl):

Area = (pi/4) / (2pi ) * 4 * pi * (3/2)^2 = 1/2 * pi * 9/4 = 9/8 pi inches ^2

OpenStudy (perl):

because surface area of sphere is 4*pi*r^2

OpenStudy (anonymous):

I got comfused on the (pi/4)/(2pi)

OpenStudy (anonymous):

Is it really 9/4

OpenStudy (perl):